# Characterisations of Sobolev spaces and constant functions over metric spaces

### v2

@@ 2.A. Lipschitz functions and the Poincaré inequality
As indicated, this was proved in [33, Theorem 1.0.1], but under a definition of the Poincaré inequality as in Lemma 2.4(2) (see [33, Definition 2.2.1]). By Lemma 2.4, this is equivalent to any of the other definitions listed in the lemma, and in particular to the version of Lemma 2.4(1) that was used in our Definition 2.2. ∎
Various examples of spaces supporting a Poincaré inequality can be found in [2], [27, Section 14.2], and [32, page 274].
@@
It is only through the application of Theorem 2.6 that the assumption of completeness enters into our considerations. An interested reader can verify that each occurrence of the assumption of “a complete space with a $(1,p)$-Poincaré inequality” could be replaced by assuming “a $(1,p-\varepsilon)$-Poincaré inequality for some $\varepsilon>0$” throughout the paper.
@@ 2.B. Hajłasz gradients and the space $\dot{M}^{1,p}(\mu)$
In contrast to the straighforward definition of the Besov norm $\dot{B}^{s}_{p,p}(\lambda)$ in (1.4), giving a meaningful definition of first order smoothness spaces over a metric measure space $(X,\rho,\mu)$ requires some more effort. The most elementary definition is probably that of Hajłasz [22]:
@@
and $\nabla f$ is the distributional gradient. This shows that [19, Theorem 1] falls under the scope of Theorem 1.6—except for details of the limit relation (1.9), which takes a sharper form in $\mathbb{R}^{d}$.
@@
Under the assumptions of Theorem 1.6, the Hajłasz–Sobolev space $M^{1,p}(\mu)$ is equivalent so several other established notions of Sobolev spaces over a metric measure space, including the Cheeger Sobolev space $\textit{Ch}^{1,p}$ of [13], the Korevaar–Schoen Sobolev space $\textit{KS}^{1,p}$ of [34], the Newtonian Sobolev space $N^{1,p}$ of [46], and the Poincaré Sobolev space $P^{1,p}$ of [23], as well as various Sobolev spaces $H^{1,p}$ associated with given systems of vector fields, which is often the preferred definition in concrete situations. See [27, Theorem 10.5.3] for $M^{1,p}=P^{1,p}=\textit{KS}^{1,p}=N^{1,p}=\textit{Ch}^{1,p}$, and [18, Theorem 10] for $H^{1,p}=P^{1,p}$ under the assumptions of Theorem 1.6 on $(X,\rho,\mu)$, and quite general assumptions on the vector fields defining $H^{1,p}$ in the last equality. The said results are formulated for the inhomogeneous versions of these spaces (involving the size of both the function and its “gradient”), but they can be readily adapted to the homogeneous versions relevant to Theorem 1.6. (Note also that, by Theorem 2.6, the assumptions of Theorem 1.6 imply those of [27, Theorem 10.5.3].)
@@
Any Heisenberg group $\mathbb{H}^{n}$, and more generally any Carnot group $\mathbb{G}$ equipped with the Carnot–Carathéodory metric $d_{cc}$ and Haar measure $\mu$, is a doubling metric measure space supporting the $(1,1)$-Poincaré inequality. Moreover, the Newtonian Sobolev space $N^{1,p}(\mathbb{G})$ (and thus the Hajłasz Sobolev space $M^{1,p}(\mathbb{G})$ by Remark 2.11) is identified with the horizontal Sobolev space $W^{1,p}_{H}(\mathbb{G})$; see [27, Section 14.2] for details and additional pointers to the literature. This shows that [38, Theorem A.3] falls under the scope of Theorem 1.6.
@@ 3. The macroscopic Poincaré inequality
For $t\approx r$ to be specified, we consider an approximate identity $\Phi_{t}$ at scale $t$ and estimate
@@ 4. How to recognise constants
With the macroscopic Poincaré inequality in our toolbox, we are in a position to give a relatively quick proof of the characterisation of constants via Besov-type conditions as in (1.5), and indeed the following slightly stronger version:
@@
then $f$ is equal to a constant almost everywhere. In particular,
$\dot{B}^{1}_{p,p}(\mu)=\{\text{constants}\}.$
@@
In the context of characterising the Schatten class properties of commutators (see Section 1.B.2 for a short discussion and [29] for the full story), yet another variant of the Besov norm naturally appears, namely
$\|f\|_{\dot{B}^{p}(\mu)}:=\Big[\iint_{X\times X}\frac{|f(x)-f(y)|^{p}}{V(x,y)^{2}}\,\mathrm{d}\mu(x)\,\mathrm{d}\mu(y)\Big]^{\frac{1}{p}}.$
Under the additional assumption of lower dimension $d$ as in (1.2) (which turns out to be natural in view of the applications in [29]), we also obtain a constancy criterion in terms of these norms as a quick corollary:
@@
Let $1\leq p\leq d<\infty$ and let $(X,\rho,\mu)$ be a doubling metric measure space having lower dimension $d$ in the sense of (1.2) and supporting the $(1,p)$-Poincaré inequality. If $f\in L^{1}_{\operatorname{loc}}(\mu)$ satisfies
@@
then $f$ is equal to a constant almost everywhere. In particular,
$\dot{B}^{p}(\mu)=\{\text{constants}\}.$
@@
Fix a ball $B_{0}:=B(x_{0},R)$. For all $x,y\in B_{0}$, we have
$\begin{split}\frac{\mu(B_{0})}{V(x,y)}&\approx\frac{V(x,R)}{V(x,y)}\quad\text{by doubling}\\ &\gtrsim\Big(\frac{R}{\rho(x,y)}\Big)^{d}\quad\text{by lower dimension}\\ &\gtrsim\Big(\frac{R}{\rho(x,y)}\Big)^{p}\quad\text{by }d\geq p\text{ and }\frac{R}{\rho(x,y)}\gtrsim 1.\end{split}$
hence
$\frac{1}{\rho(x,y)^{p}}\lesssim\frac{1}{V(x,y)}\frac{\mu(B_{0})}{R^{p}},$
@@
$\frac{|f(x)-f(y)|^{p}}{\rho(x,y)^{p}V(x,y)}\lesssim\frac{|f(x)-f(y)|^{p}}{V(x,y)^{2}}\frac{\mu(B_{0})}{R^{p}}.$
Under the assumption (4.4), the right-hand side is integrable over $B_{0}\times B_{0}$, and hence is the left-hand side. Since this is true for every ball $B_{0}$, we have verified assumption (4.2) of Proposition 4.1, and the said proposition shows that $f$ is constant. ∎
@@ 5. The lower bound for the Sobolev norm
Our aim in the rest of the paper is to prove the characterisation of the Hajłasz–Sobolev space formulated Theorem 1.6, our main result. We begin with the easy direction. This short section is a relatively straightforward adaptation of [19, Section 2]. Moreover, this direction is valid under much more general assumptions than the full Theorem 1.6.
@@ 6. Maximal versus mean oscillations
In contrast to the easy argument of Section 5, the upper bound of the Sobolev norm is much more delicate and requires some preparations. We note that the Euclidean argument of [19, Section 3] and its Carnot group extension [38, Appendix A] proceed by first considering test functions of class $C^{1,1}$ (i.e., $f\in C^{1}$ whose gradient $\nabla f$ is a Lipschitz mapping), and using Taylor expansions with second-order error terms in the estimates. Since these are not available in the present generality, this section develops some alternative tools that have no obvious counterpart in the previous literature.
@@
A deeper result is the estimate in the converse direction; this is the main result of this section. A similar bound, but with the $\liminf$ in both $\operatorname{lip}f$ and $g_{f}$ replaced by $\limsup$, is known [32, Proposition 4.3.3]; however, a $\liminf$ on the right-hand side is essential for us in view of subsequent estimates in Lemma 8.1.
@@ 7. Lipschitz approximation in $L^{p,\infty}(\nu_{p})$
Recall that the main result of the previous section, Proposition 6.3, was proved for locally Lipschitz functions only. In order to prove Theorem 1.6 without such a priori assumptions, we will need to make use of some approximation results. The density of Lipschitz functions in Sobolev spaces is well known: see [27, Theorem 8.2.1] for their density in another Sobolev-type space $N^{1,p}$, and [27, Corollary 10.2.9] for the coincidence $M^{1,p}=N^{1,p}$, both under the same assumptions that we are using. However, we need Lipschitz approximation in the a priori larger (by Proposition 5.1) space of functions with $\|m_{f}\|_{L^{p,\infty}(\nu_{p})}<\infty$, and such results do not seem to be available in the previous literature, except in the special cases of $\mathbb{R}^{d}$ [19] and Carnot groups [38], where they are relatively easy consequences of standard mollification by convolution.
@@
Thus, the oscillation of the mollification $\Phi_{t}f$ are controlled by an average of oscillations of the original $f$ at the same scale $r$. This is our substitute for the Euclidean estimate [19, proof of Lemma 3.2]
@@
This is our analogue of [19, Lemma 3.2], using Lemma 7.3 as a substitute of (7.4). In the proof of [19, Lemma 3.2], using a proper norm equivalent to the quasi-norm $\|\ \|_{L^{p,\infty}}$ and the translation invariance of the Lebesgue measure, it is immediate that
@@
This is the strong type bound corresponding to (7.6). The required (7.6) then follows by choosing some $1\leq q_{0}<q<q_{1}\leq\infty$ and interpolating (7.7) between these two values. See e.g. [28, Theorem 2.2.3] for a version covering the case at hand, with an $L^{q,\infty}$-to-$L^{q,\infty}$-bound (7.6) as a conclusion. ∎

### v3

@@ 2.A. Lipschitz functions and the Poincaré inequality
As indicated, this was proved in [38, Theorem 1.0.1], but under a definition of the Poincaré inequality as in Lemma 2.6(2) (see [38, Definition 2.2.1]). By Lemma 2.6, this is equivalent to any of the other definitions listed in the lemma, and in particular to the version of Lemma 2.6(1) that was used in our Definition 2.4. ∎
Various examples of spaces supporting a Poincaré inequality can be found in [3], [32, Section 14.2], and [37, page 274].
@@
It is only through the application of Theorem 2.8 that the assumption of completeness enters into our considerations. An interested reader can verify that each occurrence of the assumption of “a complete space with a $(1,p)$-Poincaré inequality” could be replaced by assuming “a $(1,p-\varepsilon)$-Poincaré inequality for some $\varepsilon>0$” throughout the paper. Note that this only concerns Theorem 1.1 on the characterisation of the Sobolev space, while Theorem 1.5 on the characterisation of constants is already independent of completeness as stated.
@@ 2.B. Hajłasz gradients and the space $\dot{M}^{1,p}(\mu)$
While several meaningful notions of a first-order Sobolev space over a metric measure space are known by now, the most straightforward definition is probably that of Hajłasz [27]:
@@
and $\nabla f$ is the distributional gradient. This shows that [23, Theorem 1] falls under the scope of Theorem 1.1—except for details of the limit relation (1.4), which takes a sharper form in $\mathbb{R}^{d}$.
@@
Under the assumptions of Theorem 1.1, the Hajłasz–Sobolev space $M^{1,p}(\mu)$ is equivalent so several other established notions of Sobolev spaces over a metric measure space, including the Cheeger Sobolev space $\textit{Ch}^{1,p}$ of [14], the Korevaar–Schoen Sobolev space $\textit{KS}^{1,p}$ of [39], the Newtonian Sobolev space $N^{1,p}$ of [51], and the Poincaré Sobolev space $P^{1,p}$ of [28], as well as various Sobolev spaces $H^{1,p}$ associated with given systems of vector fields, which is often the preferred definition in concrete situations. See [32, Theorem 10.5.3] for $M^{1,p}=P^{1,p}=\textit{KS}^{1,p}=N^{1,p}=\textit{Ch}^{1,p}$, and [22, Theorem 10] for $H^{1,p}=P^{1,p}$ under the assumptions of Theorem 1.1 on $(X,\rho,\mu)$, and quite general assumptions on the vector fields defining $H^{1,p}$ in the last equality. The said results are formulated for the inhomogeneous versions of these spaces (involving the size of both the function and its “gradient”), but they can be readily adapted to the homogeneous versions relevant to Theorem 1.1. (Note also that, by Theorem 2.8, the assumptions of Theorem 1.1 imply those of [32, Theorem 10.5.3].)
@@
Any Heisenberg group $\mathbb{H}^{n}$, and more generally any Carnot group $\mathbb{G}$ equipped with the Carnot–Carathéodory metric $d_{cc}$ and Haar measure $\mu$, is a doubling metric measure space supporting the $(1,1)$-Poincaré inequality. Moreover, the Newtonian Sobolev space $N^{1,p}(\mathbb{G})$ (and thus the Hajłasz Sobolev space $M^{1,p}(\mathbb{G})$ by Remark 2.13) is identified with the horizontal Sobolev space $W^{1,p}_{H}(\mathbb{G})$; see [32, Section 14.2] for details and additional pointers to the literature. This shows that [43, Theorem A.3] falls under the scope of Theorem 1.1.
@@ 3. The macroscopic Poincaré inequality
For $t\sim r$ to be specified, we consider an approximate identity $\Phi_{t}$ at scale $t$ and estimate
@@ 4. How to recognise constants
With the macroscopic Poincaré inequality in our toolbox, we are in a position to give a relatively quick proof of the characterisation of constants via Besov-type conditions as in Theorem 1.5. We have the following slightly stronger version of case (1) of the said theorem:
@@
then $f$ is equal to a constant almost everywhere. In particular, the Besov space defined in Theorem 1.5 satisfies
$\dot{B}^{1}_{p,p}(\mu)=\{\operatorname{constants}\}.$
@@
Under the additional assumption of lower dimension $d$ as in Definition 2.1 we also obtain the following variant of the constancy criterion, which turns out to be natural in view of the commutator applications in [34].
@@
Let $1\leq d<\infty$ and let $(X,\rho,\mu)$ be a doubling metric measure space having lower dimension $d$ and supporting the $(1,d)$-Poincaré inequality. If $0<p\leq d$ and $f\in L^{p}_{\operatorname{loc}}(\mu)$ satisfies
@@
then $f$ is equal to a constant almost everywhere. In particular, the Besov space defined in Theorem 1.5 satisfies
$\dot{B}^{p}(\mu)=\{\operatorname{constants}\}\qquad\forall\ p\in(0,d].$
@@
Case $p=d$. Fix a ball $B_{0}:=B(x_{0},R)$. For all $x,y\in B_{0}$, we have
$\frac{\mu(B_{0})}{V(x,y)}\sim\frac{V(x,R)}{V(x,y)}\gtrsim\Big(\frac{R}{\rho(x,y)}\Big)^{d}$
by doubling in the first step and lower dimension $d$ in the second one. Hence
$\frac{1}{\rho(x,y)^{d}}\lesssim\frac{1}{V(x,y)}\frac{\mu(B_{0})}{R^{d}},$
@@
$\frac{|f(x)-f(y)|^{d}}{\rho(x,y)^{d}V(x,y)}\lesssim\frac{|f(x)-f(y)|^{d}}{V(x,y)^{2}}\frac{\mu(B_{0})}{R^{d}}.$
Under the assumption (4.4), the right-hand side is integrable over $B_{0}\times B_{0}$, and hence is the left-hand side. Since this is true for every ball $B_{0}$, we have verified assumption (4.2) of Proposition 4.1 with $p=d$, and the said proposition shows that $f$ is constant.
Case $p\in(0,d)$. We reduce this to the previous case with the help of the following elementary observation. For $z\in\mathbb{C}$, or more generally an element of a normed space, we define its truncation at level $K>0$ by
$z_{K}:=\begin{cases}z,&\text{if }|z|\leq K,\\ \displaystyle K\frac{z}{|z|},&\text{else}.\end{cases}$ (4.5)
Then $|z_{K}|\leq K$ (obviously), and it is easy to check that $|a_{K}-b_{K}|\leq 2|a-b|$ for any two elements $a,b$. Hence
$|a_{K}-b_{K}|^{d}=|a_{K}-b_{K}|^{d-p}|a_{K}-b_{K}|^{p}\leq(2K)^{d-p}(2|a-b|)^{p}=2^{d}K^{d-p}|a-b|^{p}.$
Given $f\in L^{p}_{\operatorname{loc}}(\mu)$, we consider the truncated functions $f_{K}$, applying (4.5) pointwise. If $f$ satisfies (4.4) for some $p\in(0,d)$, then $f_{K}\in L^{\infty}_{\operatorname{loc}}(\mu)\subseteq L^{d}_{\operatorname{loc}}(\mu)$ satisfies
$\frac{|f_{K}(x)-f_{K}(y)|^{d}}{V(x,y)^{2}}\leq 2^{d}K^{d-p}\frac{|f(x)-f(y)|^{p}}{V(x,y)^{2}}\in L^{1}_{\operatorname{loc}}(\mu\times\mu).$
By the case $p=d$ that we already proved, it follows that $f_{K}$ is equal to some constant $c_{K}$ almost everywhere. Since $f_{K}=f=f_{K^{\prime}}$ on $\{f\leq\max(K,K^{\prime})\}$, it follows that $c_{K}=c_{K^{\prime}}=:c$ is independent of $K$ as soon as $K$ is large enough so that $\{f\leq K\}$ has positive measure. Since $f_{K}\to f$ pointwise as $K\to\infty$, it follows that $f=c$ almost everywhere. ∎
We record the following corollary in a specific measure space, improving a recent result of [20, Proposition 6.4]; see Remark 4.7 below for comparison and motivation.
4.6 Corollary.
Let $n\geq 0$ and $\lambda>0$, and consider the measure $\mathrm{d}m_{\lambda}(x)=x_{n+1}^{2\lambda}\,\mathrm{d}x$ on $\mathbb{R}^{n+1}_{+}$. Then
$\dot{B}^{p}(m_{\lambda})=\{\operatorname{constants}\}\qquad\forall\ p\in(0,n+1].$
Proof.
By [34, Proposition 4.2], $(\mathbb{R}^{n+1}_{+},|x-y|,m_{\lambda})$ is a doubling metric measure space of lower dimension $n+1$ and supporting the $(1,1)$ (and hence in particular the $(1,n+1)$) Poincaré inequality. Hence Corollary 4.6 is a specialisation of Corollary 4.3, with $d=n+1$, to this concrete setting. ∎
4.7 Remark.
The metric measure space $(\mathbb{R}^{n+1},|x-y|,m_{\lambda})$ of Corollary 4.6 is the so-called “Bessel setting” arising from the analysis of the Bessel differential operator
$\Delta_{\lambda}:=-\sum_{j=1}^{n+1}\frac{\partial^{2}}{\partial x_{j}^{2}}-\frac{2\lambda}{x_{n+1}}\frac{\partial}{\partial x_{n+1}}.$
As a tool for related commutator estimates, [20, Proposition 6.4] recently proved that
$\dot{B}^{p}(m_{\lambda})\cap C^{2}(\mathbb{R}_{+}^{n+1})=\{\operatorname{constants}\}\qquad\forall\ p\in(0,n+1].$ (4.8)
Corollary 4.6 shows that the restriction to $C^{2}$ functions can be removed, illustrating the power of our abstract methods, even when specialised to a concrete case like the Bessel setting.
This is used for commutator estimates as follows. For so-called Bessel–Riesz transforms $R_{\lambda,j}:=\partial_{j}\Delta_{\lambda}^{-\frac{1}{2}}$, [20, Proposition 5.2] proves that
$\|b\|_{\dot{B}^{p}(m_{\lambda})}\lesssim\|[b,R_{\lambda,j}]\|_{S^{p}(L^{2}(m_{\lambda}))}.$
In combination with (4.8), this shows that $b,R_{\lambda,j}]\in S^{p}(L^{2}(m_{\lambda}))$ if and only if $b$ is constant—under the a priori assumption that $b\in C^{2}(\mathbb{R}^{n+1}_{+})$. With Corollary 4.6 in place of (4.8), this a priori assumption may be dropped. Using the abstract Corollary 4.3 instead, this reasoning is extended to more general domains $(X,\rho,\mu)$ and operators $T$ in place of $(\mathbb{R}_{+}^{n+1},|x-y|,m_{\lambda})$ and $R_{\lambda,j}$ in [34].
For these commutator applications, one only needs case $p=d>1$ of Corollary 4.3 (which is all that we had in an earlier version of this paper). Namely, if $b,T]\in S^{p}(L^{2}(\mu))$ for some $p\in(0,d]$ and $d>1$, then we can use the elementary inclusion $S^{p}\subseteq S^{q}$ for $p\leq q$ of the Schatten classes to conclude that $b,T]\in S^{d}(L^{2}(\mu))$ as well. Then it suffices to know that $\|b\|_{\dot{B}^{d}(\mu)}\lesssim\|[b,T]\|_{S^{d}(\mu)}$ (as is proved under quite general assumptions in [34]; this needs, however, $d>1$ with strict inequality) and that $\dot{B}^{d}(\mu)=\{\operatorname{constants}\}$ (guaranteed by case $p=d$ of Corollary 4.3; this is valid for $d\geq 1$, with equality allowed) to conclude that $b=\operatorname{constant}$.
@@ 5. The lower bound for the Sobolev norm
Our aim in the rest of the paper is to prove the characterisation of the Hajłasz–Sobolev space formulated Theorem 1.1, our main result. We begin with the easy direction. This short section is a relatively straightforward adaptation of [23, Section 2]. Moreover, this direction is valid under much more general assumptions than the full Theorem 1.1.
@@ 6. Maximal versus mean oscillations
In contrast to the easy argument of Section 5, the upper bound of the Sobolev norm is much more delicate and requires some preparations. We note that the Euclidean argument of [23, Section 3] and its Carnot group extension [43, Appendix A] proceed by first considering test functions of class $C^{1,1}$ (i.e., $f\in C^{1}$ whose gradient $\nabla f$ is a Lipschitz mapping), and using Taylor expansions with second-order error terms in the estimates. Since these are not available in the present generality, this section develops some alternative tools that have no obvious counterpart in the previous literature.
@@
A deeper result is the estimate in the converse direction; this is the main result of this section. A similar bound, but with the $\liminf$ in both $\operatorname{lip}f$ and $g_{f}$ replaced by $\limsup$, is known [37, Proposition 4.3.3]; however, a $\liminf$ on the right-hand side is essential for us in view of subsequent estimates in Lemma 8.1.
@@ 7. Lipschitz approximation in $L^{p,\infty}(\nu_{p})$
Recall that the main result of the previous section, Proposition 6.3, was proved for locally Lipschitz functions only. In order to prove Theorem 1.1 without such a priori assumptions, we will need to make use of some approximation results. The density of Lipschitz functions in Sobolev spaces is well known: see [32, Theorem 8.2.1] for their density in another Sobolev-type space $N^{1,p}$, and [32, Corollary 10.2.9] for the coincidence $M^{1,p}=N^{1,p}$, both under the same assumptions that we are using. However, we need Lipschitz approximation in the a priori larger (by Proposition 5.1) space of functions with $\|m_{f}\|_{L^{p,\infty}(\nu_{p})}<\infty$, and such results do not seem to be available in the previous literature, except in the special cases of $\mathbb{R}^{d}$ [23] and Carnot groups [43], where they are relatively easy consequences of standard mollification by convolution.
@@
Thus, the oscillation of the mollification $\Phi_{t}f$ are controlled by an average of oscillations of the original $f$ at the same scale $r$. This is our substitute for the Euclidean estimate [23, proof of Lemma 3.2]
@@
This is our analogue of [23, Lemma 3.2], using Lemma 7.3 as a substitute of (7.4). In the proof of [23, Lemma 3.2], using a proper norm equivalent to the quasi-norm $\|\ \|_{L^{p,\infty}}$ and the translation invariance of the Lebesgue measure, it is immediate that
@@
This is the strong type bound corresponding to (7.6). The required (7.6) then follows by choosing some $1\leq q_{0}<q<q_{1}\leq\infty$ and interpolating (7.7) between these two values. See e.g. [33, Theorem 2.2.3] for a version covering the case at hand, with an $L^{q,\infty}$-to-$L^{q,\infty}$-bound (7.6) as a conclusion. ∎